# Decommissioning and post-irradiation examination of the LHC beam dumps

### v1

@@ Abstract
The LHC beam dumps are responsible for the safe absorption of the Large Hadron Collider (LHC) particle beams. In 2018, the two 6.4-tonne beam dumps that had been in operation since the LHC’s startup in 2008 were removed and replaced with upgraded versions. Endoscopic inspections of these beam dumps and experimental high-intensity proton-beam irradiation of material samples raised concerns about the structural integrity of the carbon-based materials in their cores. It was therefore decided to undertake an accelerated project of dismantling and post-irradiation examination of the removed dumps as part of a wider program of work to ensure the safe operation of the LHC beam dumps in the coming years. This paper describes the decommissioning process for the two beam dumps carried out at CERN between 2021 and 2023, covering the preparatory studies, practical challenges encountered, and solutions implemented. It details the establishment of an operational framework, including the preparation of the working environment, the development of a method for cutting the irradiated -thick duplex stainless-steel vessel, and the cut sequencing. Additionally, the paper presents the findings derived from the post-irradiation examination of the different carbon-based core materials subjected to deposited energy densities up to . The extruded graphite plates within the vessel exhibited a cracking pattern, which was likely due to the dynamic response of the device upon beam impact, and their retaining rings were found to be displaced. Despite minor signs of surface deterioration, the expanded graphite sheets were intact, and the isostatic graphite blocks showed no evidence of material degradation. These findings were consistent across both beam dumps. Finally, this manuscript discusses the implications of these results for the safe operation of the beam dumps during LHC Run 3 (2022–2025) and for the design and construction of Run 3 operational spares.
@@ II Beam-induced energy deposition and response of the dump to beam impact
Upon beam impact, a secondary particle shower develops along the length of the dump due to nuclear interactions between the particle beam and the carbon-based materials. The energy deposited in the dump core materials can be computed using the Monte Carlo-based FLUKA simulation package [16, 17, 18], as shown in Figs. 4(b)–4(d). Figure 4(a) shows the maximum transverse energy density as a function of the longitudinal position within the dump across Run 1 and Run 2.
Although the LBDS was designed and verified for a high level of dependability (rated SIL4 according to IEC 61508 [19]), partial dilution failures cannot be entirely excluded [20], and these can result in an increase in the peak energy densities achieved in the dump. Three dilution failures with dumped energy exceeding occurred during the 9-year operation of the dumps [21]. The most critical of these, with a dumped energy of , resulted in energy densities in the core comparable to those produced during a full-intensity dump () in nominal operation during Run 2. A full summary of the beam parameters for different LHC runs is presented in Table 1.
@@ IV.3 Operative radiation protection
Figure 10 shows a schematic of the beam dump dismantling worksite. This was enclosed by concrete blocks measuring in length, in thickness, and in height. These blocks served to reduce the radiation background in the surrounding area and to comply with the building’s area classification requirements, i.e., Supervised Radiation Area [ ] [26]. The worksite was divided into different subareas: the main working area (working area 1, covering ), where the main cutting activities were performed; an area dedicated to the RW preconditioning (working area 2, covering ); a storage bunker for RW, completely shielded to reduce the ambient dose equivalent rate in the two main working areas; and two airlocks, one dedicated to personnel and one for materials. The worksite area delimited by the shielding blocks was classified as a Limited Stay Radiation Area, with a maximum allowed dose rate of [26].
@@ VII.2 Isostatic graphite
Similar to the LD expanded graphite sheets, the surface of the isostatic graphite blocks displayed a distinct pattern of small black dots, indicating the path of the beam. This pattern eased the determination and marking of extraction positions across different energy densities, enabling a comparative assessment. The final coring positions, overlaid with the corresponding energy density map, are shown in Fig. 30. Due to the HD blocks being shrink-fitted, the presence of the activated vessel was unavoidable, requiring the extraction to be performed by a robot with a coring tool, as shown in Fig. 31. A vacuum cleaner was installed around the coring tool to minimize the spread of particles and potential contamination during the operation.
@@
SEM images were taken to assess any surface changes between the corings; however, no significant differences were observed. Figure 33 shows the results of a chemical analysis performed using EDS [44]. The darkened areas showed the presence of carbon (C), oxygen (O), and nitrogen (N), in line with the observations for the expanded graphite sheets. Hardness tests were carried out following the Shore D standard, as the material was nonmetallic and too porous for conventional micro- or macro-hardness testing. The hardness of all corings fell within the range 72–85, with no significant correlation to the energy density to which each coring was subjected.
@@ VIII.2 Design of Run 3 spare beam dumps
In response to the observed cracking of the extruded graphite plates, the plates for the Run 3 operational spares have been replaced by carbon-fiber-reinforced carbon (CFRC) plates. These materials, which have already been used in beam-intercepting devices exposed to high levels of energy deposition, are attractive due to their significantly greater strength and enhanced resistance to crack initiation and propagation. The retaining plates now feature larger outgassing holes, which are intended to reduce local stress concentrations and decrease the likelihood of crack initiation. This modification aims to ensure proper containment and compression of the LD expanded graphite sheets. In addition, a thorough assessment of the retaining rings has been conducted, resulting in a complete redesign to improve their stability and the implementation of effective measures for securing them firmly within the dumps. Furthermore, the design has been optimized for disassembly by eliminating the tack welds between the vessel and the retaining ring, instead employing a pin-lock configuration that secures the retaining ring within the vessel’s groove.

### v2

@@ Abstract
The LHC beam dumps are responsible for the safe absorption of the Large Hadron Collider (LHC) particle beams. In 2018, the two 6.4-tonne beam dumps that had been in operation since the LHC’s startup in 2008 were removed and replaced with upgraded versions. Endoscopic inspections of these beam dumps and experimental high-intensity proton-beam irradiation of material samples raised concerns about the structural integrity of the carbon-based materials in their cores. It was therefore decided to undertake an accelerated project of dismantling and post-irradiation examination of the removed dumps as part of a wider program of work to ensure the safe operation of the LHC beam dumps in the coming years. This paper describes the decommissioning process for the two beam dumps carried out at CERN between 2021 and 2023, covering the preparatory studies, practical challenges encountered, and solutions implemented. It details the establishment of an operational framework, including the preparation of the working environment, the development of a method for cutting the irradiated -thick duplex stainless-steel vessel, and the cut sequencing. Additionally, the paper presents the findings derived from the post-irradiation examination of the different carbon-based core materials subjected to deposited energy densities up to . The extruded graphite plates within the vessel exhibited a cracking pattern, which was likely due to the dynamic response of the device upon beam impact, and their retaining rings were found to be displaced. Despite minor signs of surface deterioration, the expanded graphite sheets were intact, and the isostatic graphite blocks showed no evidence of material degradation. These findings were consistent across both beam dumps. Finally, this manuscript discusses the implications of these results for the safe operation of the beam dumps during LHC Run 3 (2022–2026) and for the design and construction of Run 3 operational spares.
@@ II Beam-induced energy deposition and response of the dump to beam impact
Upon beam impact, a secondary particle shower develops along the length of the dump due to nuclear interactions between the particle beam and the carbon-based materials.
A crucial component of the LBDS is the set of four horizontal and six vertical kicker magnets, known as dilution kickers, located immediately downstream of the LHC extraction kicker magnets. These magnets apply a time-dependent transverse kick to the beam bunches, creating a sweeping pattern across the front face of the beam dump. This pattern evolves laterally along the extraction tunnel, increasing the transverse separation between bunches and thereby reducing the peak energy density in the dump core [3].
The energy deposited in the dump core materials can be computed using the Monte Carlo-based FLUKA simulation package [16, 17, 18], as shown in Figs. 4(b)–4(d). Figure 4(a) shows the maximum transverse energy density as a function of the longitudinal position within the dump across Run 1 and Run 2. The sweeping pattern, visible in these figures, originates at the center of the transverse plane and progresses in a counter-clockwise direction.
Although the LBDS was designed and verified to meet a high level of dependability (rated SIL4 according to IEC 61508 [19]), partial dilution failures, where some of the dilution kicker magnets fail to operate, cannot be entirely excluded [20]. Such failures can lead to reduced lateral bunch separation, resulting in higher peak energy densities in the beam dump.
Three dilution failures with dumped energy exceeding occurred during the 9-year operation of the dumps [21]. The most critical of these, with a dumped energy of , resulted in energy densities in the core comparable to those produced during a full-intensity dump () in nominal operation during Run 2. A full summary of the beam parameters for different LHC runs is presented in Table 1.
@@ IV.3 Operative radiation protection
Figure 10 shows a schematic of the beam dump dismantling worksite. This was enclosed by concrete blocks measuring in length, in thickness, and in height. These blocks served to reduce the radiation background in the surrounding area and to comply with the building’s area classification requirements, i.e., Supervised Radiation Area [ ] [26]. Here, denotes the dose at a depth of in tissue, and is used to assess radiation exposure in controlled areas.
The worksite was divided into different subareas: the main working area (working area 1, covering ), where the main cutting activities were performed; an area dedicated to the RW preconditioning (working area 2, covering ); a storage bunker for RW, completely shielded to reduce the ambient dose equivalent rate in the two main working areas; and two airlocks, one dedicated to personnel and one for materials. The worksite area delimited by the shielding blocks was classified as a Limited Stay Radiation Area, with a maximum allowed dose rate of [26].
@@ VII.2 Isostatic graphite
Similar to the LD expanded graphite sheets, the surface of the isostatic graphite blocks displayed a distinct pattern of small black dots, indicating the path of the beam. This pattern eased the determination and marking of extraction positions across different energy densities, enabling a comparative assessment. Control samples were also taken outside the sweep pattern to evaluate the sensitivity of the material to beam impact in less-irradiated areas. The final coring positions, overlaid with the corresponding energy density map, are shown in Fig. 30. Due to the HD blocks being shrink-fitted, the presence of the activated vessel was unavoidable, requiring the extraction to be performed by a robot with a coring tool, as shown in Fig. 31. A vacuum cleaner was installed around the coring tool to minimize the spread of particles and potential contamination during the operation.
@@
SEM images were taken to assess any surface changes between the corings; however, no significant differences were observed. Figure 33 shows the results of a chemical analysis performed using EDS [44]. The darkened areas showed the presence of carbon (C), oxygen (O), and nitrogen (N), in line with the observations for the expanded graphite sheets. The hardness of all core samples ranged from 72–85 according to the Shore D standard. This standard was used as the material was nonmetallic and too porous for conventional micro- or macro-hardness testing. The hardness tests showed no clear dependence on extraction location. Furthermore, no significant correlation was observed between hardness and the energy density each sample experienced along the beam sweep.
@@ VIII.2 Design of Run 3 spare beam dumps
In response to the observed cracking of the extruded graphite plates, the plates for the Run 3 operational spares have been replaced by carbon-fiber-reinforced carbon (CFRC) plates. These materials, which have already been used in beam-intercepting devices exposed to high levels of energy deposition, are attractive due to their significantly greater strength and enhanced resistance to crack initiation and propagation [47]. The retaining plates now feature larger outgassing holes, which are intended to reduce local stress concentrations and decrease the likelihood of crack initiation. This modification aims to ensure proper containment and compression of the LD expanded graphite sheets. In addition, a thorough assessment of the retaining rings has been conducted, resulting in a complete redesign to improve their stability and the implementation of effective measures for securing them firmly within the dumps. Furthermore, the design has been optimized for disassembly by eliminating the tack welds between the vessel and the retaining ring, instead employing a pin-lock configuration that secures the retaining ring within the vessel’s groove.